OpenStudy (anonymous):

A certain type of wter lily spreads to double the area it covers every year ten yes after being planted in our lake those water lilies cover a quarter of it's surface how many more years will it be before they cover the entire surface of the lake? Plz explain I don't get this at all

OpenStudy (anonymous):

Every year, the area of the lilies double. So say one year, they cover 1 square metre. The following year, they cover double that area: 2 square metres. The next year, they cover double that: 4 square metres... The progression goes: 1 2 4 8 16 32 ... \[n^{\text{years passed}}\] So if you're starting with an area of \(\frac{1}{4}\), what will be the area the following year?

OpenStudy (anonymous):

um 1/5

OpenStudy (anonymous):

1/5 is unfortunately not double 1/4. Maybe if I write it in equation form: \[2 \times \frac{1}{4}=?\]

OpenStudy (anonymous):

That's what you get when you add them, not when you multiply them.

OpenStudy (anonymous):

When you multiply 2\(\times\)3, what you're saying is "Add 3 twice", or "3+3". So when you say \(2 \times \frac{1}{4}\), what you mean is "Add \frac{1}{4} twice", or "\(\frac{1}{4}+\frac{1}{4}\)" \[\]